# Large Language Models in Argument Mining: A Survey

### v2

@@ 4 Modelling Paradigms in the LLM Era
With the advent of Large Language Models (LLMs), the landscape of argument mining has undergone a significant shift. Traditional pipeline-based or shallow neural architectures are increasingly being complemented—or replaced—by foundation models that exhibit strong generalization, reasoning, and few-shot learning capabilities. This chapter categorizes the emerging modelling paradigms into four key directions: prompt-based adaptation, fine-tuning and instruction-tuning, retrieval-augmented generation, and synthetic supervision.
@@ 4.1 Prompt-based Argument Mining
Prompting has become a dominant paradigm in adapting LLMs for argument mining, enabling zero-shot and few-shot performance without task-specific supervision. Many recent works demonstrate the effectiveness of simple or chain-of-thought prompts for various sub-tasks. For example, in stance detection, Zhao et al. (2024) introduced ZeroStance, a framework that uses ChatGPT to both generate data and perform inference in a zero-shot setting. Similarly, Ma et al. (2024b) proposed Chain-of-Stance, a prompting scheme to induce intermediate reasoning steps for improved stance inference.
Prompting has also been used to assess the sufficiency or quality of arguments. Li et al. (2025) explore prompt-based counterfactual calibration strategies to mitigate stance-related biases in LLMs. Chen and Eger (2025) use prompts for manipulation checks in evaluating the effect of emotional content on argument convincingness.
These studies underscore prompting’s flexibility, interpretability, and rapid adaptability across domains and tasks.
4.2 Instruction Tuning and Multi-task Fine-tuning
While prompting alone can offer promising zero-shot performance, several works have proposed fine-tuning LLMs using argument-specific instructions or multi-task objectives. Li et al. (2024b) introduced a unified framework trained on multi-task objectives covering key point generation, evidence detection, and argument quality estimation. Their model demonstrated strong generalization across argument summarization and evaluation benchmarks.
In the legal domain, instruction-tuned models such as ArgLegalSumm (Elaraby and Litman, 2022) improve summarization by aligning model outputs with argumentative structures. Elaraby et al. (2023) further enhanced legal summarization using re-ranking mechanisms driven by LLM-computed argument salience scores.
In stance detection, models such as LOGIC (Lee et al., 2024) demonstrate that LLM-generated rationales can guide small student models through intermediate supervision—blending instruction tuning with model distillation.
4.3 Retrieval-Augmented and Knowledge-Grounded Generation
Given the inherent limitations of even the largest LLMs, retrieval-based models have emerged as a hybrid solution for argument mining. These approaches augment the model’s context window by retrieving relevant evidence, past arguments, or external knowledge.
For instance, OpenDebateEvidence (Roush et al., 2024) introduces a massive-scale retrieval-augmented dataset for claim-centric argument generation and summarization. In debate summarization, Syed et al. (2023) employ frame-oriented representations derived from retrieved background knowledge to enhance the generation of structured summaries.
Some models integrate argumentation theory into the retrieval process. Ruiz-Dolz et al. (2023) use argument graph semantics to guide the retrieval of supporting and attacking evidence, while Palmieri (2024) apply critical questions as filters during retrieval and evaluation.
4.4 Synthetic Data and LLM-generated Supervision
Synthetic supervision has emerged as a promising avenue for bootstrapping argument mining tasks at scale. LLMs can be used to synthesize high-quality training data for low-resource domains or tasks lacking annotated corpora.
Wagner et al. (2025) show that LLM-synthesized stance datasets outperform traditional few-shot fine-tuning, enabling rapid domain adaptation. Similarly, Khosravani et al. (2024) use LLM-generated key point coverage scores to construct automatic evaluation metrics for argument summarization.
This paradigm is not only cost-efficient but also enables the control and modulation of specific attributes such as diversity (van der Meer et al., 2024), exhaustiveness (Khosravani et al., 2024), and coherence (Shan and Lu, 2025)—factors crucial to high-quality argument outputs.
5 Argument Evaluation
Argument evaluation is a higher-level task in argument mining that assesses the quality, strength, and sufficiency of arguments. It encompasses multiple aspects including convincingness, sufficiency, bias, quality ranking, and interpretability. The growing research in this area has led to the development of both theoretical frameworks and empirical methods leveraging machine learning and large language models (LLMs).
5.1 Convincingness Evaluation
Convincingness refers to how persuasive or compelling an argument is. Early studies such as Chalaguine and Schulz (2017) showed that simple linguistic and structural features could suffice for assessing convincingness in online debates. Later, more advanced models were introduced. Gu et al. (2018) explored incorporating topic aspects in evaluating comment convincingness on social platforms, while Potash et al. (2019) proposed ranking mechanisms for argument passages.
The impact of emotions has also been investigated. Chen and Eger (2025) introduced an LLM-based manipulation check system to explore whether emotional language biases human and machine assessments of argument convincingness.
5.2 Sufficiency and Quality
Sufficiency evaluation addresses whether an argument provides enough support for a given claim. A key contribution is Gurcke et al. (2021), who assess sufficiency via conclusion generation, operationalizing the idea that sufficient arguments should enable accurate inference of the claim. Similarly, Liu et al. (2024b) proposed CASA, a causality-aware method for argument sufficiency assessment.
Further formal analysis of sufficiency was provided by Borg and Bex (2024), who explored the concepts of minimality, necessity, and sufficiency in the context of explanation and argumentation. Quality assessment tasks often intersect with sufficiency. For instance, Favreau et al. (2022) applied learning-to-rank BERT models to evaluate argument quality in educational settings.
5.3 Automatic Evaluation and End-to-End Frameworks
Recent works aim to integrate evaluation into end-to-end pipelines. Li et al. (2024b) proposed a multi-task dataset encompassing summarization and evaluation components, enabling automatic scoring of argument relevance and strength. Ruiz-Dolz et al. (2023) introduced a graph-based semantic network to evaluate debate quality using argument graph theory and neural modeling.
A novel coverage metric for summarization tasks was proposed by Khosravani et al. (2024), emphasizing exhaustiveness and accuracy in key point extraction, which reflects evaluation criteria in real-world persuasive writing. From a logic-oriented angle, Burrieza and Yuste-Ginel (2021) examined evaluation through justification logic in multi-agent environments. In education, Winkle et al. (2020) analyzed how teachers apply critical questions to evaluate student arguments, while Rach et al. (2020) benchmarked argument search and evaluation systems in dialogue settings.
5.4 Theoretical and Domain-Specific Perspectives
Beyond purely computational methods, there is growing attention to theoretical and domain-specific evaluation. Palmieri (2024) applied the Argument Mapping Technique (AMT) in corporate controversies to assess argument quality using critical questions. In legal contexts, Grundler et al. (2024) developed AMELIA, an evaluation benchmark for Italian legal documents.
Biases in argument evaluation have also been studied. Diana et al. (2020) investigated adaptive instruction methods to mitigate value-based biases. Baccini and Hartmann (2022) formalized the well-known myside bias using Bayesian models.
@@ 6 Future Directions
As argument mining transitions into the era of large language models (LLMs), a wide range of opportunities emerge—alongside challenges that must be systematically addressed. This chapter synthesizes key limitations identified across prior sections and outlines promising future research directions spanning technical, theoretical, and practical dimensions.
6.1 Model Interpretability and Controllability
While LLMs show strong performance across tasks such as stance detection, argument generation, and summarization, their outputs often lack transparency and controllability. In high-stakes domains such as law (Elaraby and Litman, 2022; Xu and Ashley, 2023), healthcare, or political discourse, black-box behavior can hinder trust and adoption.
Future work should prioritize model interpretability, perhaps via structured intermediate representations, symbolic scaffolds (e.g., argument diagrams), or rationale-based generation (Lee et al., 2024). Alignment techniques like self-refinement, chain-of-thought prompting, or rule-augmented decoding may further improve controllability in argumentative outputs.
6.2 Data Quality, Bias, and Hallucination Risks
Several studies, such as Chen and Eger (2025) and Li et al. (2025), have raised concerns regarding the influence of emotion, ideology, or prompt phrasing on argument quality and stance. LLMs trained on massive web corpora may inherit or amplify biases, leading to unbalanced or misleading arguments.
Moreover, hallucination remains a critical challenge in argument generation, particularly when models are required to retrieve or construct supporting evidence. This calls for the integration of fact verification systems, grounding in structured knowledge, or retrieval-augmented techniques (Roush et al., 2024).
6.3 Generalization and Robustness
Despite notable successes in zero- and few-shot settings, many models struggle to generalize across domains (e.g., legal, social media, political debate). Argument summarization models trained on English legal corpora often fail on dialogic or multilingual data (Zhao et al., 2023). Furthermore, domain shifts and adversarial examples can degrade performance in stance classification or evidence retrieval tasks.
Research should focus on domain adaptation, robustness to noisy or adversarial inputs, and multi-lingual evaluation—perhaps through multi-task, curriculum, or meta-learning frameworks.
6.4 Synthetic Supervision and Benchmark Quality
The rise of LLM-generated datasets (e.g., ZeroStance (Zhao et al., 2024), OpenDebateEvidence (Roush et al., 2024)) introduces a double-edged sword. While they enable rapid scaling and coverage, they risk data leakage (pretraining contamination), superficial reasoning, or poor coverage of edge cases.
Future efforts should investigate hybrid data curation pipelines that blend human oversight, synthetic diversity control, and task-specific quality metrics (Khosravani et al., 2024). Benchmarking standards need to evolve accordingly, with support for open-ended outputs, argument graphs, or quality dimensions like sufficiency, diversity, and persuasiveness.
6.5 Integration with Argumentation Theory
Despite decades of theoretical work on argumentation (e.g., Toulmin models, Dung frameworks), most LLM-based methods remain disconnected from formal theory. Argument sufficiency, support/attack structure, or minimality principles are rarely encoded into training objectives or evaluation metrics.
Bridging this gap may involve using theory-guided prompts, symbolic constraints, or structure-aware generation and evaluation systems (Borg and Bex, 2024; Palmieri, 2024). Such alignment could improve model reasoning, transparency, and downstream decision support.
6.6 Human-Centered and Interactive Argument Mining
Finally, the shift toward interactive systems—e.g., debate coaching bots, civic deliberation platforms, or educational feedback tools—demands models that support real-time, multi-turn, and user-aligned argument mining. This includes tracking dialog acts, modeling myside bias (Baccini and Hartmann, 2022), and adapting feedback to user values or learning goals (Diana et al., 2020).
Future systems must integrate HCI design, personalization strategies, and continual learning mechanisms to realize the full societal potential of LLM-based argumentation systems.
@@ 7 Conclusion
In this survey, we reviewed the evolution of argument mining in the era of large language models (LLMs), outlining the key tasks—ranging from claim detection and evidence retrieval to stance classification, argument summarization, and evaluation. We categorized modeling approaches before and after the LLM shift, highlighting the increasing role of prompt-based, generative, and retrieval-augmented techniques.
While LLMs have significantly advanced performance across argument mining tasks, they also introduce new challenges related to interpretability, robustness, data quality, and alignment with argumentation theory. As research moves forward, we emphasize the need for better grounding, human-centric interaction, and theoretically informed evaluation.
We hope this survey offers a comprehensive reference for scholars and practitioners seeking to build more reliable, transparent, and cognitively aligned systems for argument understanding and generation.

### v3

@@ 6 Modelling Paradigms in the LLM Era
Large Language Models (LLMs) have transformed argument mining from a pipeline of narrow, task-specific classifiers into a set of interaction paradigms in which one or more foundation models collaborate with retrieval modules, symbolic structures or lightweight adapters. Rather than ask whether an architecture is “better” than another, current research asks how LLMs are enlisted as zero-shot predictors, few-shot imitators, knowledge retrievers, judges, data generators, or all of the above. This section synthesises the dominant paradigms, emphasising the design rationales, empirical pay-offs, and unresolved pitfalls.
@@ 6.1 Prompt-based Argument Mining
Early demonstrations showed that even naïve declarative prompts (“Is the following sentence a claim?”) enable GPT-3 to match feature-rich baselines on component segmentation and stance detection without gradient updates (Chen et al., 2024). Subsequent work moved beyond template engineering to embed task structure directly inside the prompt. Altemeyer et al. (2025) inserted chain-of-thought (CoT) exemplars that model the logical leap between premise and conclusion; the reasoning trace, rather than the final label, accounted for most of the performance gain on sufficiency and convincingness. Prompting can also express task hierarchy: the PITA framework conditions relation prediction on previously elicited component spans within the same conversation context and reports significant improvements on Reddit and legal briefs (Sun et al., 2024b). Role prompts further reduce topical and demographic bias: framing the model as a Debate Judge or Devil’s Advocate attenuates polarity skew in political fact-checking (Dayanik and Padó, 2020).
Yet prompt brittleness persists. Small lexical tweaks can flip stance labels, and CoT traces lengthen latency in interactive writing tools. Robust prompt ensembles and self-consistency voting help but introduce computational overhead; empirical studies suggest latency grows almost linearly with ensemble size (Zhang et al., 2025). Designing prompts that capture argumentation theory while remaining compact therefore remains an open problem.
6.2 Parameter-efficient Instruction and Multi-task Tuning
Zero-shot prompting rarely suffices for domain-sensitive genres such as biomedical abstracts or statutory law. Instruction tuning—fine-tuning on a mixture of task-specific exemplars framed as natural-language instructions—bridges much of this gap. A single T5-base model (Chung et al., 2024) jointly instruct-tuned on claim segmentation, relation prediction and quality scoring yields consistent gains across all three tasks on IAM and ARIES, while adapters keep GPU memory within budget (Cheng et al., 2022; Gemechu et al., 2024). In the legal domain, ArgLegalSumm couples instruction tuning with an argument salience re-ranker, improving summary coherence in Italian case law (Elaraby et al., 2023). Cross-domain continual learning is more fragile: when a biomedical corpus is appended late in the schedule, earlier gains on social-media stance detection erode, echoing catastrophic forgetting observed in multi-task translation (Pang et al., 2024).
Recent work therefore explores parameter-efficient updates: LoRA (Hu et al., 2022), prefix-tuning and bit-fit adapters enable domain adaptation with two orders of magnitude fewer trainable parameters while preserving general AM capability (Li et al., 2024b). Still unsettled is how to balance tasks of differing label density: curriculum strategies that start with coarse component labels before introducing noisy quality scores appear promising, but systematic ablation is scarce.
6.3 Retrieval-Augmented Generation and Reasoning
LLMs remain constrained by context windows and training cut-off dates. Retrieval-augmented generation (RAG) mitigates both limits by injecting external evidence into the prompt. In debate summarisation, RAG systems retrieve topically relevant counter-arguments and background facts, reducing hallucination and increasing factual density in abstractive summaries (Syed et al., 2023). The OpenDebateEvidence benchmark pairs fifteen million statements with evidence links, enabling end-to-end retrieval-generation pipelines that outperform text-only baselines by a meaningful margin (Roush et al., 2024). Retrieval can also be argument-theoretic: semantic-graph matching over Kialo debate trees yields higher recall of undercutters and mitigations than purely lexical search (Ruiz-Dolz et al., 2023).
Several challenges loom. Evidence ranking often inherits popularity biases, favouring mainstream sources over minority viewpoints. Long evidence snippets may eclipse the claim inside the prompt, forcing heuristic truncation. Finally, retrieval latency threatens interactive applications; hybrid dense-sparse indexing and passage caching alleviate but do not solve the issue.
6.4 Synthetic Supervision and Data Augmentation
LLMs themselves can generate training data, especially where expert annotation is scarce. Wagner et al. (2025) show that LLM-synthesised stance corpora, when filtered for diversity, raise low-resource performance more than in-context few-shot learning. Similarly, LLM-generated key-point coverage scores serve as weak labels for automatic evaluation metrics, correlating well with human judgements (Khosravani et al., 2024). Synthetic supervision also facilitates controlled experiments: topic-balanced sampling can equalise stance priors across political entities, supporting fairness audits (Mehrabi et al., 2022).
The downside is distributional drift. Synthetic arguments tend to be lexically fluent, hedge-free and uniformly polite, properties that can over-regularise fine-tuned models. Provenance cards documenting prompt templates, sampling temperatures, rejection rates and post-editing guidelines are therefore being proposed as a best practice (DBLP:conf/facc/Paullada2021). Legal uncertainty over copyright of machine-generated text remains unresolved.
@@ 7 Future Directions
As argument mining transitions into the era of large language models (LLMs), a wide range of opportunities emerge alongside challenges that must be systematically addressed. This chapter synthesizes key limitations identified across prior sections and outlines promising future research directions spanning technical, theoretical, and practical dimensions.
7.1 Interpretability and Controllability
LLM outputs can be impressively fluent yet logically opaque. Black-box convincingness scores or abstractive summaries are difficult to trust in law, medicine or policy making, where explanations must be auditable and contestable. Promising directions include symbolic scaffolding, in which an LLM populates a Toulmin-style or Walton-style template before emitting a verdict (Lee et al., 2024; Borg and Bex, 2024); rationale-aligned distillation, where model predictions are accepted only if an intermediate reasoning trace satisfies argument-quality criteria (Zhao et al., 2024); and rule-augmented decoding, which filters candidate continuations against critical-question constraints during generation (Palmieri, 2024). Empirical studies show that such structured intermediates improve user trust more than generic saliency maps, yet rigorous human-in-the-loop evaluations are still rare.
7.2 Data Quality, Bias and Hallucination
LLM training corpora contain ideological skew, demographic imbalance and source-credibility noise. These artefacts surface in persuasive writing: emotionally charged but weakly supported arguments are often overrated by LLM judges (Chen and Eger, 2025). Counterfactual calibration and actor masking reduce partisan bias in stance detection (Li et al., 2025; Dayanik and Padó, 2020), but systematic audits across domains remain scarce. Hallucination is the sister problem: retrieval-augmented generation mitigates unsupported claims yet inherits popularity and language biases from search engines (Roush et al., 2024). Future benchmarks should combine bias slices e.g. position papers from under-represented groups— with factual stress tests to quantify both issues jointly.
7.3 Generalisation and Robustness
Most LLM studies report single-domain results; cross-genre transfers often collapse. Debate summarisers tuned on English legal cases lose more than a third of their content fidelity when applied to multi-party chat (Zhao et al., 2023). Adversarial paraphrase attacks flip stance predictions with minimal lexical edits (Gienapp et al., 2020). Potential remedies include curriculum-style multi-task fine-tuning, meta-learning over domain clusters, and evaluation suites that mix genres, languages and modalities within the same test fold.
7.4 Synthetic Supervision and Benchmark Integrity
LLM-generated corpora such as ZeroStance (Zhao et al., 2024) and OpenDebateEvidence (Roush et al., 2024) accelerate data growth but raise three red flags: (i) contamination models may re-encounter their own pre-training shards; (ii) superficial reasoning synthetic counter-claims are stylistically rich yet logically shallow; (iii) blind spot edge cases rarely appear in temperature-controlled sampling. Hybrid pipelines that blend human editing, diversity-aware sampling and automatic quality-metrics (e.g. coverage and exhaustiveness scores from Khosravani et al. (2024)) are a promising compromise. Benchmark leaders should publish provenance cards documenting prompt templates, sampling temperatures and rejection rates.
7.5 Theory-Grounded Modelling
Most LLM pipelines treat argument quality as a monolithic scalar. Argumentation theory, however, differentiates logical sufficiency, dialectical relevance and rhetorical effectiveness. Embedding these distinctions into prompts, loss functions or evaluation metrics is still nascent. Early attempts to weave Walton’s critical questions into dialogue evaluation improved error localisation but slowed inference (Palmieri, 2024). Future work could explore neuro-symbolic hybrids that use graph-based logic checkers as fast filters before LLM re-drafting.
7.6 Human-Centred and Interactive AM
Educational feedback tools, civic debate platforms and policy consultation interfaces require models that operate with users, not merely for them. That includes tracking my-side bias (Baccini and Hartmann, 2022), adapting critique to a learner’s knowledge state (Diana et al., 2020), and providing actionable, value-aligned suggestions. Reinforcement learning from argumentative feedback, real-time uncertainty estimates and personalisation via retrieval of prior user arguments are promising building blocks.
@@ 8 Conclusion
The field of Argument Mining has come a long way, moving from narrow, feature-engineered pipelines trained on small, task-specific corpora to a rich ecosystem powered by large language models. Today’s systems can identify claims, map support and attack relations, gauge stance, assess quality, and even summarize key points, all within a unified framework that leverages prompting, few-shot adaptation, retrieval of background evidence, and synthetic data generation. Alongside these gains, however, have come new challenges: the opacity of model reasoning, the risk of bias and hallucination, brittleness under domain shift, and questions of provenance and ethical deployment.
Looking ahead, progress will hinge on our ability to reconcile raw performance with transparency, robustness and theoretical grounding. This means developing interpretable abstractions e.g. symbolic scaffolds, rationale-based decoding and uncertainty-aware metrics, that expose the “why” behind LLM decisions; constructing datasets and benchmarks that capture the full spectrum of argument modalities, languages and genres; auditing for demographic, ideological and stylistic bias; and designing interactive, human-centric systems that guide rather than replace expert judgment. By weaving together advances in model technology, data design and evaluation practice, we can build argument-mining tools that are not only accurate, but also trustworthy, equitable and broadly useful in the real world.